OpenStudy (anonymous):

Given that BD Perpendicular to AC , what is the length of ? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):



OpenStudy (anonymous):

the problem says this:Given that BD ACWhat is the length of AC?Reply Using Drawing

OpenStudy (anonymous):



OpenStudy (anonymous):

When you have a right triangle you can find a missing side using the Pythagorean Theorem: a^2 + b^2 = c^2

OpenStudy (anonymous):

so will you walk me through it step by step please!?

OpenStudy (anonymous):

Yeah(:

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Okay so we know a and b are the legs of the triangle (the sides that make up the ninety degree angle, so in this case your line segments BA and AC.) Your c value would be your hypotenuse which is your line BC. When you put your values into the Pythagorean Theorem it would look like this: 9.4^2 + b^2 = 23.2^2 where your b value is the length of AC Using simple algebra rewrite as 23.3^2 -9.4^2 = b^2 542.89 - 18.8 = b^2 524.09 = b^2 So, take the sqrt of that and you get?

OpenStudy (anonymous):

sqrt of 454.53= 21.31970919 sqrt of 294376.1101= 542.5643834

OpenStudy (anonymous):

hold up.... Where did you get those values from?

OpenStudy (anonymous):

you mean take sqrt of 524.09?

OpenStudy (anonymous):

You only need to find the length of AC hun, so yeah... The sqrt of 524.09 is the answer

OpenStudy (anonymous):

oh...lol ok. thanks

OpenStudy (anonymous):

No problem jabbers(:

OpenStudy (anonymous):

22.89?

OpenStudy (anonymous):

Correct!:D

OpenStudy (anonymous):

thanks for helping me

OpenStudy (anonymous):

it's my pleasure to be of help :)

OpenStudy (anonymous):

will u help with another 1 if if i put in new question box?